# Impact chronology of leftover planetesimals

### v1

@@ 3.2 Impact chronology and population decline
where the constants are weights adding up to 1, the variables are e-folding times, and e is the base of the natural logarithm; the number of terms required turned out to be usually four and sometimes three. I prefer this formulation rather than using a stretched exponential – – because the multi-exponential formulation yields a better fit, and the fitting constants yield insight about the initial conditions and fundamental dynamical processes; specifically the number of terms reveal the number of sinks or processes involved. The results of the fits are given in Table 1. For impacts on the Earth the values of and are rather similar across the models, being identical to within 25% and having values near 35 Myr and 100 Myr; the difference increase to to 37% for .
@@
For Mars the middle two time constants in the ABI model are similar to those for the Earth, but they carry different weights; the martian impact time constants are more different from Earth’s for the other two models. For the fraction of remaining planetesimals, the same equation also works well. In particular, for the GT model the first three time constants are similar to impacts on Earth, while the weights are identical within factors of 25%. For Mars the time constants of its impact chronology for the DD model resembles the overall decline of that population. This indicates that the impact rate is modulated by the overall rate of decline, or removal, of the planetesimals from the reservoirs.
@@
These results beg the question whether the coefficients are related to the initial fraction of Earth (and Mars) crossers. In the GT, ABI and DD models the initial fraction of Earth crossers is 70%, 50% and 35%, which are close to the coefficients for impacts on Earth. For the ABI and DD models for Earth, so the remaining particles not controlled by Earth and only by Mars impact the Earth on a timescale . For impacts on Mars only three terms are needed, and for the ABI and DD models martian impacts occur mostly on timescales and , in contrast to for the GT model. For all models the impact rate on the terrestrial planets is related to the overall decline of the whole population As such, if no planetesimals remain to impact Mars because most of them are gone after a few times , then the rate of martian impacts must also decline in tandem.
@@
In summary, the time constants for both the impacts and the remaining fraction seem to be clustered in four groups: one group at Myr, another at Myr, a third at Myr and a fourth that is generally Myr. This clustering suggests that the rate of impacts is partially set by the rate of decline of planetesimals, which in turn is set by their leakage rate from their initial reservoirs. For comparison: Nesvorný et al. (2017) used three terms and found Myr, Myr and Myr for the rate of impacts on the Earth, even though their study only included the main asteroid belt plus E-belt subjected to an episode of giant planet migration. These similar values support my earlier hypothesis that the time constants are features of the system. Furthermore, the crater chronologies of Neukum et al. (2001) and Werner et al. (2023) have e-folding timescales of Myr, which is of the same order as , indicating that the lunar craters mostly reflect the impacts from an ancient Mars-crossing population. This result implies that the number of Earth crossers at the time that the Moon formed had to be low, probably less than half that of the Mars crossers.
3.2.1 The evolution of the perihelion distribution
@@
In Fig. 6 I plot the evolution of the cumulative perihelion distribution for each model in the left column, and the complementary cumulative distributions of the impacts on Earth (blue) and Mars (red), as well as the fraction of planetesimals remaining (black) and the total planetesimals that are removed (orange) in the right column. The GT model is the top row, the middle is for ABI, and the bottom for DD. In all models the cumulative distribution rapidly changes shape, with the fraction of Earth crossers dropping quickly, and the fraction of Mars crossers reducing more slowly. At later stages, usually after 700 Myr, almost all of the remaining planetesimals have au. The rate of change of the distribution can be approximately quantified as follows. For all three models the fraction of Earth crossers normalised by the remaining population decreases exponentially with an e-folding time of Myr, while for the Mars crossers in the DD model the fraction decreases with an e-folding time of 220 Myr while it is increased to 430 Myr for the GT and ABI models. Thus, for the DD model the fraction of Earth crossers normalised by the remaining population declines at the same rate as that of the Mars crossers, while for the ABI and GT models the Mars crossing fraction declines much slower.
@@
I argued that the initial perihelion distribution is the main constraint in determining the impact chronology on the Earth and Mars as well as global population decline. This argument is evident from the figure: the initial population decline and impact rate for the GT model is much more rapid than for the ABI and DD models, which is evident from the fitting constants in Table 1. For the ABI and GT models the terrestrial impacts track the global decline for the first 100 Myr and begin to deviate from there, with the normalised rate of impacts on Mars being always slower than on Earth. In contrast, the terrestrial impact chronology closely tracks all the planetesimals depletion in the GT and ABI models, while Mars’ impact chronology is initially off to a slower start but its impact rate converges after 200 Myr. Beyond this point, in all models the impact chronology is set by the dynamical evolution of the Mars crossers, which evolve on a timescale of about 80–100 Myr (Michel et al., 2000), in other words , because encounters with Mars will need to supply the planetesimals to Earth, either directly or via the .
@@
The behaviour of the remaining fraction of planetesimals is different: it levels off after a few hundred Myr of evolution due to planetesimals being parked on long-lived orbits beyond Mars. Some of these slowly leak onto Mars-crossing orbits due to chaotic diffusion (Morbidelli & Nesvorný, 1999), and some of these planetesimals collide with the terrestrial planets.
@@ 3.3 Removal mechanism behind the fitted timescales
A greater insight into the origin of the timescales is to investigate which planetesimals are removed at what time, and from where. Figure 7 is a heat map showing the original semi-major axis and eccentricity of all planetesimals that are removed from the simulation at different time quartiles. The top left panel shows the removal for the first quartile of the population, the top right panel the second quartile, the bottom left shows the third quartile, and the bottom right shows the last quartile. The top four squares are for the GT model, and the bottom four are for the ABI model. Throughout the simulation planetesimals are removed from all over, as evidenced by the blue dots, and instead I focus on the orange regions.
@@
For the GT model in the top left panel, which lasts for 7 Myr and is comparable to , removed planetesimals initially reside on Earth-crossing orbits with semi-major axis between 1.2 and 2.5 au. There are clear indicators of rapid removal of planetesimals beginning near 2 au and 2.5 au even at low initial eccentricity, corresponding to the and 3:1. The next quartile, shown in the top right panel, occurs between 7 Myr and 21 Myr and still has the densest part on Earth-crossing orbits but at shorter semi-major axis, between 1.1 au and 1.6 au. The high density region shifts to shorter semi-major axis in the third panel, lasting from 21 Myr to 52 Myr. Combined, these quartiles last for about 1.5. The reason for this shift towards shorter semi-major axis is that it takes longer to eliminate these objects through collision or by scattering them into the or the 3:1. As such, it appears likely that is related to the rate of removal of Earth crossers. The last panel, lasting 948 Myr, shows the slowest-declining population, which are Mars crossers with semi-major axis shorter than 2 au. This last quartile is dominated by and – although the latter has a low coefficient – and suggests that is the timescale of evolution of Mars crossers.
@@
For the ABI model the top left panel has the highest density of planetesimals initially on Earth-crossing orbits with initial semi-major axes between 1.3 au and 2 au, with also some bodies in the – see the vertical points near 2 au at low eccentricity. This quartile lasts for 12 Myr, comparable to , suggesting that is related to removal through the and high-eccentricity Earth (and Venus) crossers. The next quartile, lasting from 12 Myr to 36 Myr, still has a large number of Earth crossers, but there are also a fair amount of Mars crossers in there; the region is still being emptied out. By now half of the planetesimals are removed. During the third quartile, lasting from 36 Myr to 91 Myr, most of the planetesimals that are removed are Mars crossers with initial semi-major axes between 1.2 au and 2 au. The rate of population decline is beginning to slow down, with the terms with Myr and Myr setting the rate, once again suggesting that is related to the removal of Earth crossers and to that of Mars crossers. The fourth quartile lasts for 900 Myr, is modulated by the terms with and at late stages, and consists mostly of Mars crossers. However, the E-belt region beyond Mars, with initial semi-major axis between 1.6 au and 2 au and perihelion au (Bottke et al., 2012), also plays a role; this region appears to play almost no role in the GT simulations.
@@
In summary, I argue that is likely a measure of the decline of high-eccentricity Earth-crossers and bodies initially in the and 3:1, describes Earth and Venus crossers, and pertains to the Mars crossers, and possibly the E-belt. The term , which is the last few percent, is due to the decline of bodies originally beyond Mars on more stable orbits such as the inner main belt.
@@
where is the encounter velocity scaled by the planet’s orbital velocity, and is the -component of this velocity – which is essentially the radial component of the encounter velocity. For the initial conditions considered in the simulations the value of for the Earth ranges from 1.5 to 2.2, and . The radius at which a collision occurs can be computed from energy and angular momentum conservation, and is given by ; when for the Earth and 1.13 for Venus. This leads to a probability of collision with the Earth or Venus of per orbit of a planetesimal that crosses the Earth’s or Venus’ orbit. The corresponding e-folding time for collision is then Myr using a typical semi-major axis of au. This timescale is far longer than the timescales for removal numerically computed here. In the simulations only of planetesimals are removed through collision with an inner planet; the rest are removed by ’ejection’, that is, they are being scattered into the secular resonance or 3:1 mean-motion resonance through repeated encounters with the inner planets. These resonances pump up the eccentricities of the planetesimals and they ultimately collide with the Sun or encounter Jupiter (Gladman et al., 1997).
To calculate the removal timescale I need an estimate of the cumulative change in semi-major axis due to scattering. For small deflections Arnold (1965) computes that the cumulative effect of distant encounters beyond logarithmically increases the change in energy through encounters closer than , resulting in an enhancement by a factor of 4.5. Since is large and deflections are small, I can extrapolate this relation for all encounters beyond , which increases it to 6.9. Therefore, a rough timescale of ’ejection’ of a planetesimal through repeated Earth encounters is about seven times shorter than that of removal through collision, resulting in an e-folding time of 60 Myr. By harmonically adding the effect of elimination through Venus encounters (Arnold, 1965) I arrive at an e-folding timescale of removal through combined Earth and Venus encounters of 30 Myr, which is comparable to found by my dynamical simulations.
@@
For Mars and so that . Elimination through ’ejection’ and collisions with Mars results in an e-folding time of about 200 Myr when au, which is twice as long as found numerically here as well as in the literature (Michel et al., 2000; Fernández & Helal, 2023), but still of the same order. Objects near Mars undergo large changes in eccentricity due to the proximity of the secular resonance, which can bring them to Earth-crossing orbits which assists in their removal. Still, I argue that is associated with the removal of planetesimals by interaction with Mars.
@@
For both models the rate of decline during the last quartile, which lasts Myr for all models, is dominated by Mars crossers, while during the second and third quartile the rate is still influenced by remaining Earth crossers. However, for the GT model Myr for the decline of the population, akin to the timescale found by Michel et al. (2000) for the Mars crossers. Given the similarity between the initial cumulative semi-major axis distribution in the ABI and DD models and the slower impact rate of the latter, it stands to reason that the initial perihelion distribution is the most important parameter controlling the impact rate. The fitting constants therefore probably represent the (initial) perihelion and semi-major axis distribution.
@@
One final point of note pertains to the total fraction remaining after 1 Gyr. For both the GT and ABI models this is about 1%, but for the DD model this is about 10%. The last value is problematic because the current mass in the main belt is about , which probably declined by about 50%-65% over the past 4 Gyr (Minton & Malhotra, 2010; Nesvorný et al., 2017; Brasser et al., 2020), with even less than that in the inner main belt. In the DD model the remaining 10% resides in the inner main belt, implying that the initial population in planetesimals at 4.5 Ga was no greater than . This is insufficient to account for the lunar craters and its highly siderophile element abundances in its mantle (Morbidelli et al., 2018; Brasser et al., 2020). Therefore I will have to discount the initial conditions of the DD model as being viable for late accretion.
@@ 4 Spherule layer constraints on the Archean impact rate and source population
It has been suggested that terrestrial spherule beds are a record of ancient terrestrial impacts (Simonson & Glass, 2004). Johnson & Melosh (2012) used the spherule layer thickness from the recorded list of known spherule beds in Glass & Simonson (2012) to compute the diameter range of the impactor that created the spherule bed. It is not clear whether this list is complete, but I will treat it as such. In Table 2 I have listed the names and ages of the largest of these spherule beds, the average impactor diameter from the range listed in Johnson & Melosh (2012), and the corresponding number of impacts on the Earth for objects with diameter km using a size-frequency distribution slope of . For completeness I added the Sudbury layer with the proposed impactor diameter using the formula from Johnson & Melosh (2012) but I do not use it in the analysis. There is also the Vredefort crater in South Africa with an approximate age of 2 Ga, which may have left traces of a bed as far as Russia (Huber et al., 2014; Allen et al., 2022). No spherule beds have been found between 0.6 and 1.7 Ga ago, but this time interval has not been extensively explored for impact spherules (Simonson & Glass, 2004).
Bottke et al. (2012) used the presence of the spherule layers and their thickness to compute the impact rate onto the Earth from a declining population of asteroids placed just beyond the orbit of Mars. Here I will use a slightly different approach. From Table 2 I compute that there were 17k impacts on Earth between 3.47 Ga and 2 Ga with diameter km. This results in an average impact rate of 11.5 Myr-1. The uncertainties are factors of a few due to the possible range of the impactor diameters. What reservoir provided these planetesimals?
@@
During the Archaean and later the cratering rate on the Earth can be computed analogous to the outer Solar System, and is simply (Brasser et al., 2025), where is the flux of objects on Earth-crossing orbits, and is the impact probability with the Earth. The results in Fig. 6 seem to apply that this relation is justified by the similar shape of the terrestrial and martian impact curves with those of removed planetesimals. Furthermore, at late times ( Myr of simulation time) most objects are entirely beyond Mars so that the impact rate is determined by the rate at which these objects become Mars crossers. This situation is similar to that of Kuiper Belt and Scattered Disc objects that leak into the realm of the giant planets from beyond Neptune (Duncan & Levison, 1997).
@@
In practice the flux is equal to (Duncan et al., 1995), where is the total number of objects in all reservoirs with km. I follow Duncan et al. (1995) and compute the rate of decline of planetesimals as the number of planetesimals that left the system () divided by the final number of the particles at the end (), and the time interval (). In other words
@@
It is worth mentioning that generally depends on the time range over which it is calculated because the rate of decline is generally not constant with time; instead, the rate of decline is a function of time itself: it decreases with increasing time. An exception is an exponential decay of the form e-t/τ, for which we have , which is constant. From the simulation output at late times . In addition, apart from the small fraction of objects that are ejected due to encounters with Jupiter, the far majority are lost due to collisions with a planet or the Sun; on the way to the Sun they are temporary Earth crossers, so that to a good approximation is equal to the rate of decline of the total remaining population (see also Fig. 6).
@@
In the simulations, the impact probability with Earth ranges between 7.3% (ABI and DD) and 9% (GT), so I will take 7.5% as the average. This impact probability implies that 227k objects with km were lost between 3.47 Ga and 2 Ga from the various small body reservoirs in the inner Solar System. Assuming that and Gyr then the aforementioned 227k objects represent a decline of 63% by 2 Ga in the population remained at 3.47 Ga. A similar value of Gyr was also found by Nesvorný et al. (2017). If instead Myr, as suggested by the impacts on the Earth and Mars, then the population remaining at 3.47 Ga would have declined a further 99.3% by 2 Ga. Such an extreme depletion can be ruled out because it would imply an implausibly large mass in the initial population at 4.5 Ga. Using Gyr implies that there were about 617k objects with km at 3.47 Ga available to supply the terrestrial impactors. The main belt between 2.1 au and 3.3 au barely declines (Minton & Malhotra, 2010; Nesvorný et al., 2017; Brasser et al., 2020), so that these 617k planetesimals must have been in the leftover population, or in the E-belt, or both.

### v3

@@ 3.2 Impact chronology and population decline
where the constants are weights adding up to 1, the variables are e-folding times, and e is the base of the natural logarithm; the number of terms required turned out to be usually four and sometimes three. I prefer this formulation rather than using a stretched exponential – – because the multi-exponential formulation yields a better fit, and the fitting constants yield insights into the initial conditions and fundamental dynamical processes; specifically, the number of terms reveals the number of sinks or processes involved. The results of the fits are given in Table 1. For impacts on the Earth the values of and are rather similar across the models, being identical to within 25% and having values near 35 Myr and 100 Myr; the difference increase to to 37% for .
@@
For impacts on Mars, the middle two time constants in the ABI model are similar to those for the Earth, but they carry different weights; the martian impact time constants are more different from Earth’s for the other two models. I found that a fit with three terms generally worked better than one with four terms. The same equation also works well for the decline of the whole population, although the last term has an e-folding time that is much longer than the simulation time. I find that the exponential fit at large , which corresponds to the term with , is only an approximation: experimentation shows that here a linear fit would also suffice, akin to the constant cratering flux assumption (Neukum et al., 1975). In particular, for the GT model the first three time constants fitting the decline of the whole population are similar to impacts on Earth, while the weights are identical within factors of 25%. For impacts on Mars the time constants of its impact chronology for the DD model resembles the overall decline of the whole population. This indicates that, to first order, the impact rate is modulated by the overall rate of decline, or removal, of the planetesimals from the reservoirs.
@@
These results beg the question whether the coefficients are related to the initial fraction of Earth (and Mars) crossers. In the GT, ABI, and DD models the initial fraction of Earth-crossers is 70%, 50% and 35%, which are close to the coefficients for impacts on Earth. For the ABI and DD models for Earth, so the remaining particles not controlled by Earth and only by Mars impact the Earth on a timescale . For impacts on Mars only three terms are needed, and for the ABI and DD models martian impacts occur mostly on timescales and , in contrast to for the GT model. For all models the impact rate on the terrestrial planets is related to the overall decline of the whole population As such, if no planetesimals remain to impact Mars because most of them are gone after a few times , then the rate of martian impacts must also decline in tandem.
@@
In summary, the time constants for both the impacts and the remaining fraction seem to be clustered in four groups: one group at Myr, another at Myr, a third at Myr and a fourth that is generally Myr. This clustering suggests that the rate of impacts is partially set by the rate of decline of planetesimals, which in turn is set by their leakage rate from their initial reservoirs. For comparison: Nesvorný et al. (2017) used three terms and found Myr, Myr and Myr for the rate of impacts on the Earth, even though their study only included the main asteroid belt plus E-belt subjected to an episode of giant planet migration. These similar values support my earlier hypothesis that the time constants are features of the system. Furthermore, the crater chronologies of Neukum et al. (2001) and Werner et al. (2023) have e-folding timescales of Myr, which is of the same order as , indicating that the lunar craters mostly reflect the impacts from an ancient Mars-crossing population. This result implies that the number of Earth-crossers at the time that the Moon formed had to be low, probably less than half that of the Mars-crossers.
3.3 The evolution of the perihelion distribution
@@
In Fig. 6 I plot the evolution of the cumulative perihelion distribution for each model and the complementary cumulative distributions of the impacts on Earth and Mars, as well as the fraction of planetesimals remaining and the total planetesimals that are removed. In all models the cumulative distribution is normalised to the number of remaining planetesimals and rapidly changes shape, with the fraction of Earth-crossers dropping quickly, and the fraction of Mars-crossers reducing more slowly. At late stages, usually after 700 Myr, almost all of the remaining planetesimals have au. The rate of change of the cumulative distribution at specific perihelion values can be approximately quantified as follows.
I denote by the fraction of remaining planetesimals that is still Earth or Mars crossing. In other words, corresponds to the cumulative value of the perihelion distribution plotted in Fig. 6. From the simulation output I conclude that for all three models the fraction of Earth-crossers – – decreases exponentially with an e-folding time of Myr. For example, after 500 Myr in the GT model the fraction of Earth-crossers while initially it is 0.7. The approximate e-folding time of the decline of the remaining population that is still Earth crossing is then Myr. For the fraction of Mars-crossers – – I find an e-folding time of 220 Myr in the DD model, while the e-folding time is increased to 430 Myr for the GT and ABI models. Thus, for the DD model the value of declines at the same rate as , but in the ABI and GT models the fraction of planetesimals that are still Mars crossing declines much more slowly. The rate of decline of ) for the Earth and Mars-crossers is generally much lower than the decline in impacts and that of the whole population.
@@
I argued that the initial perihelion distribution is the main constraint in determining the impact chronology on the Earth and Mars as well as global population decline. This argument is evident from the figure: the initial population decline and impact rate for the GT model is much more rapid than for the ABI and DD models, which is evident from the fitting constants in Table 1. For the ABI and GT models, the terrestrial impacts track the global decline for the first 100 Myr and begin to deviate from there, with the normalised rate of impacts on Mars being always lower than on Earth. In contrast, the terrestrial impact chronology closely tracks all the planetesimals depletion in the GT and ABI models, while Mars’ impact chronology is initially off to a slower start but its impact rate converges after 200 Myr. Beyond this point, in all models the impact chronology is set by the dynamical evolution of the Mars-crossers, which evolve on a timescale of about 80–100 Myr (Michel et al., 2000), in other words , because encounters with Mars will need to supply the planetesimals to Earth, either directly or via the .
@@
The behaviour of the remaining fraction of planetesimals is different: after a few hundred million years of evolution it levels off due to planetesimals being parked on long-lived orbits beyond Mars. Some of these slowly leak onto Mars-crossing orbits due to chaotic diffusion (Morbidelli & Nesvorný, 1999), and some of these planetesimals collide with the terrestrial planets.
@@ 3.4 Removal mechanism behind the fitted timescales
To gain a greater insight into the origin of the timescales, I investigated which planetesimals are removed at what time, and from where. Figure 7 is a heat map showing the original semi-major axis and eccentricity of all planetesimals that are removed from the simulation at different time quartiles. The top left panel shows the removal for the first quartile of the population, the top right panel the second quartile, the bottom left shows the third quartile, and the bottom right shows the last quartile. The top four squares are for the GT model, and the bottom four are for the ABI model. Throughout the simulation planetesimals are removed from all over, as evidenced by the blue dots, and instead I focus on the orange regions.
@@
For the GT model in the top left panel, which lasts for 7 Myr and is comparable to , removed planetesimals initially reside on Earth-crossing orbits with semi-major axis between 1.2 and 2.5 au. There are clear indicators of rapid removal of planetesimals beginning near 2 au and 2.5 au even at low initial eccentricity, corresponding to the and 3:1. The next quartile, shown in the top right panel, occurs between 7 Myr and 21 Myr and still has the densest part on Earth-crossing orbits but at shorter semi-major axis, between 1.1 au and 1.6 au. The high density region shifts to shorter semi-major axis in the third panel, lasting from 21 Myr to 52 Myr. Combined, these quartiles last for about 1.5. The reason for this shift towards shorter semi-major axis is that it takes longer to eliminate these objects through collision or by scattering them into the or the 3:1. As such, it appears likely that is related to the rate of removal of Earth-crossers. The last panel, lasting 948 Myr, shows the most slowly declining population, which are Mars-crossers with semi-major axes between 1.5 au and 2 au. This last quartile is dominated by and – although the latter has a low coefficient – which suggests that is the timescale of evolution of Mars-crossers.
@@
For the ABI model the top left panel has the highest density of planetesimals initially on Earth-crossing orbits with initial semi-major axes between 1.3 au and 2 au, with also some bodies in the – see the vertical points near 2 au at low eccentricity. This quartile lasts for 12 Myr, comparable to , suggesting that is related to removal through the and high-eccentricity Earth (and Venus) crossers. The next quartile, lasting from 12 Myr to 36 Myr, still has a large number of Earth-crossers, but there are also a fair amount of Mars-crossers in there; the region is still being emptied out. By now half of the planetesimals are removed. During the third quartile, lasting from 36 Myr to 91 Myr, most of the planetesimals that are removed are Mars-crossers with initial semi-major axes between 1.2 au and 2 au. The rate of population decline begins to decrease, with the terms with Myr and Myr setting the rate, once again suggesting that is related to the removal of Earth-crossers and to that of Mars-crossers. The fourth quartile lasts for 900 Myr, is modulated by the terms with and at late stages, and consists mostly of Mars-crossers. However, the E-belt region beyond Mars, with initial semi-major axis between 1.6 au and 2 au and perihelion au (Bottke et al., 2012), also plays a role; this region appears to play almost no role in the GT simulations.
@@
In summary, I argue that is likely a measure of the decline of high-eccentricity Earth-crossers and bodies initially in the and 3:1, describes Earth and Venus crossers, and pertains to the Mars-crossers, and possibly the E-belt. The term , which is the last few percent, is due to the decline of bodies originally beyond Mars on more stable orbits such as the inner main belt.
@@
where is the encounter velocity scaled by the planet’s orbital velocity, and is the -component of this velocity – which is essentially the radial component of the encounter velocity. For the initial conditions considered in the simulations the value of for the Earth ranges from 1.5 to 2.2, and . The radius at which a collision occurs can be computed from energy and angular momentum conservation, and is given by ; when I compute that for the Earth and 1.13 for Venus. This leads to a probability of collision with the Earth or Venus of per orbit of a planetesimal that crosses the Earth’s and/or Venus’s orbit. The corresponding e-folding time for collision is then Myr using a typical semi-major axis of au. This timescale is far longer than the timescales for removal numerically computed here. In the simulations only of planetesimals are removed through collision with an inner planet; the rest are removed by ‘ejection’, that is, they are being scattered into the secular resonance or 3:1 mean-motion resonance through repeated encounters with the inner planets. These resonances pump up the eccentricities of the planetesimals and they ultimately collide with the Sun or encounter Jupiter (Gladman et al., 1997).
To calculate the removal timescale I need an estimate of the cumulative change in semi-major axis due to scattering. I wrote a Fortran program that implements the Monte Carlo method of Arnold (1965) using Öpik’s equations as shown in Valsecchi et al. (1997, 2003) for encounters with the Earth and Venus; the addition of Mars is trivial but was not considered. Sinks include collisions with Venus and Earth, a perihelion beyond 4.5 au, entrance into the secular resonance, and ejection (. I considered only encounters within 60 . The initial semi-major axis, perihelion distance and inclination were drawn from the Rayleigh and logistic distributions discussed in Sect. 2. The mean dynamical lifetime is Myr (2) and the median is 36.1 Myr. The lifetime distribution is roughly logistic in , with fitting parameters (corresponding to Myr), and width . It is known that Öpik’s method does not adequately describe the decline of planetesimals (Dones et al., 1999), so I consider these estimates to be upper limits. In any case, a rough estimate of the average dynamical lifetime of a planetesimal through repeated Earth and Venus encounters is 70 Myr, which is about 2.
@@
For Mars and so that . Elimination through encounters and collisions with Mars results in an e-folding time of about 200 Myr when au, which is twice as long as found numerically here as well as in the literature (Michel et al., 2000; Fernández & Helal, 2023), but still of the same order. Objects near Mars undergo large changes in eccentricity due to the proximity of the secular resonance; this can send them onto Earth-crossing orbits, which in turn leads to their rapid removal. Still, I argue that is associated with the removal of planetesimals by interaction with Mars.
@@
For both models the rate of decline during the last quartile, which lasts Myr for all models, is dominated by Mars-crossers, while during the second and third quartile the rate is still influenced by remaining Earth-crossers. However, for the GT model Myr for the decline of the population, akin to the timescale found by Michel et al. (2000) for the Mars-crossers. Given the similarity between the initial cumulative semi-major axis distribution in the ABI and DD models and the lower impact rate of the latter, it stands to reason that the initial perihelion distribution is the most important parameter controlling the impact rate. The fitting constants therefore probably represent the (initial) perihelion and semi-major axis distribution.
@@
One final point of note pertains to the total fraction remaining after 1 Gyr. For both the GT and ABI models this is about 1%, but for the DD model this is about 10%. The last value is problematic because the current mass in the main belt is about , which probably declined by about 50%-65% over the past 4 Gyr (Minton & Malhotra, 2010; Nesvorný et al., 2017; Brasser et al., 2020), with even less than that in the inner main belt. In the DD model the remaining 10% resides in the inner main belt, implying that the initial population in planetesimals at 4.5 Ga was no greater than . This is insufficient to account for the lunar craters and the abundances of highly siderophile elements in the lunar mantle (Morbidelli et al., 2018; Brasser et al., 2020). Therefore, given the simulations that I have run here, I will have to cautiously discount the initial conditions of the DD model as being viable for late accretion.
@@ 4 Spherule layer constraints on the Archean impact rate and source population
It has been suggested that terrestrial spherule beds are a record of ancient terrestrial impacts (Simonson & Glass, 2004). Johnson & Melosh (2012) used the spherule layer thickness from the recorded list of known spherule beds in Glass & Simonson (2012) to compute the diameter range of the impactor that created the spherule bed. It is not clear whether this list is complete, but I will treat it as such. In Table 2 I have listed the names and ages of the largest of these spherule beds, the average impactor diameter from the range listed in Johnson & Melosh (2012), and the corresponding number of impacts on the Earth for objects with diameter km using a size-frequency distribution slope of . For completeness I added the Sudbury layer with the proposed impactor diameter using the formula from Johnson & Melosh (2012), but I do not use it in the analysis. There is also the Vredefort crater in South Africa with an approximate age of 2 Ga, which may have left traces of a bed as far as Russia (Huber et al., 2014; Allen et al., 2022). No spherule beds have been found between 0.6 and 1.7 Ga ago, but this time interval has not been extensively explored for impact spherules (Simonson & Glass, 2004).
Bottke et al. (2012) used the presence of the spherule layers and their thickness to compute the impact rate onto the Earth from a declining population of asteroids placed just beyond the orbit of Mars. Here I will use a slightly different approach. From Table 2 I compute that there were 17k impacts on Earth between 3.47 Ga and 2 Ga of planetesimals with diameter km. This results in an average impact rate of 11.5 Myr-1. The uncertainties are factors of a few due to the possible range of the impactor diameters. What reservoir provided these planetesimals?
@@
During the Archean and later, the cratering rate on the Earth can be computed analogous to the outer Solar System, and is simply (Brasser et al., 2025), where is the flux of objects on Earth-crossing orbits, and is the impact probability with the Earth. The results in Fig. 6 seem to apply that this relation is justified by the similar shape of the terrestrial and martian impact curves with those of removed planetesimals. Furthermore, at late times ( Myr of simulation time) most objects are entirely beyond Mars so that the impact rate is determined by the rate at which these objects become Mars-crossers. This situation is similar to that of Kuiper belt and scattered disc objects that leak into the realm of the giant planets from beyond Neptune (Duncan & Levison, 1997).
@@
In practice, the flux of objects leaking in from beyond Mars is equal to (Duncan et al., 1995), where is the total number of objects in all reservoirs with km. I followed Duncan et al. (1995) and computed the rate of decline of planetesimals as the number of planetesimals that left the system () divided by the final number of the particles at the end (), and the time interval (). In other words,
@@
It is worth mentioning that generally depends on the time range over which it is calculated because the rate of decline is generally not constant with time; instead, the rate of decline is a function of time itself: it decreases with increasing time. An exception is an exponential decay of the form e-t/τ, for which we have , which is constant. From the simulation output at late times . In addition, apart from the small fraction of objects that are ejected due to encounters with Jupiter, the far majority are lost due to collisions with a planet or the Sun; on the way to the Sun they are temporary Earth-crossers, so that to a good approximation is equal to the rate of decline of the total remaining population (see also Fig. 6). However, care should be taken in using . One could argue that the relation for impacts on Earth and the population decline are not identical, violating the relation, but that is only true if one assumes that is constant with time. This is generally not the case: grows with time as the population declines and planetesimals are eliminated. Only at late stages, when the planetesimals are leaking in from beyond Mars, is the assumption that is constant approximately valid.
@@
In the simulations, the impact probability with Earth ranges between 7.3% (ABI and DD) and 9% (GT), so I will take 7.5% as the average. This impact probability implies that 227k objects with km were lost between 3.47 Ga and 2 Ga from the various small body reservoirs in the inner Solar System. Assuming that and Gyr then the aforementioned 227k objects represent a decline of 63% by 2 Ga in the population remained at 3.47 Ga. A similar value of Gyr was also found by Nesvorný et al. (2017). Using Gyr implies that there were about 617k objects with km at 3.47 Ga available to supply the terrestrial impactors. The main belt between 2.1 au and 3.3 au barely declines (Minton & Malhotra, 2010; Nesvorný et al., 2017; Brasser et al., 2020), so that these 617k planetesimals must have been in the leftover population, or in the E-belt, or both.
The fit to impacts on Earth and Mars suggest that Myr rather than Gyr from the fit to the remaining planetesimals. This would suggest a discrepancy between the two fits and that choosing Gyr is incorrect. However, upon closer inspection, this is not the case: it can be shown by doing a Taylor expansion of the fits at large (e.g. at 800 Myr) that the rates of decline of the impacts and the whole population are equal to within a factor of two. In addition, if indeed Myr, then the population remaining at 3.47 Ga would have declined a further 99.3% by 2 Ga. Such an extreme depletion can be ruled out because it would imply an implausibly large mass in the initial population at 4.5 Ga that violates constraints from the lunar cratering record and the terrestrial zircon record (Brasser et al., 2021).